# A Modified Cosmic Brane Proposal for Holographic Renyi Entropy

### v2

@@ Abstract
We propose a new formula for computing holographic Renyi entropies in the presence of multiple extremal surfaces. Our proposal is based on computing the wave function in the basis of fixed-area states and assuming a diagonal approximation for the Renyi entropy. For Renyi index $n\geq 1$, our proposal agrees with the existing cosmic brane proposal for holographic Renyi entropy. For $n<1$, however, our proposal predicts a new phase with leading order (in Newton’s constant $G$) corrections to the cosmic brane proposal, even far from entanglement phase transitions and when bulk quantum corrections are unimportant. Recast in terms of optimization over fixed-area states, the difference between the two proposals can be understood to come from the order of optimization: for $n<1$, the cosmic brane proposal is a minimax prescription whereas our proposal is a maximin prescription. We demonstrate the presence of such leading order corrections using illustrative examples. In particular, our proposal reproduces existing results in the literature for the PSSY model and high-energy eigenstates, providing a universal explanation for previously found leading order corrections to the $n<1$ Renyi entropies.
@@ 1 Introduction
The Renyi entropies at integer $n$ can be computed using the replica trick in the boundary CFT Calabrese:2004eu . The insight of Ref. Lewkowycz:2013nqa , as we shall review in Sec. 2.1, was to propose that the corresponding dominant bulk gravitational saddle is replica symmetric. Then, quotienting the bulk geometry by the replica symmetry, one obtains a geometry with an additional conical defect of opening angle $\frac{2\pi}{n}$ anchored to $\partial R$. Such a conical defect could equivalently be interpreted as being induced by the insertion of a “cosmic brane” of appropriate tension Dong:2016fnf ; Lewkowycz:2013nqa . This cosmic brane proposal provides an analytic continuation away from integer $n$, and the RT formula follows from it in the limit $n\to 1$. Moreover, it was shown in Ref. Dong:2016fnf that the refined Renyi entropy is then computed by the area of the cosmic brane in this new spacetime, providing a generalization of the RT formula. Henceforth, we will refer to either of the above proposals for Renyi entropy and refined Renyi entropy as the cosmic brane proposal.
In this paper, we will demonstrate that while the cosmic brane proposal is correct in many situations, it can fail even at leading order (in $G$) in large regions of parameter space. In particular, we will show that such corrections appear quite generically for Renyi index $n<1$ in the presence of multiple extremal surfaces. Such corrections have previously been noticed by Refs. Dong:2021oad ; Akers:2022max in the PSSY model of black hole evaporation Penington:2019kki . Moreover, the results of Ref. Murthy:2019qvb interpreted in a holographic context also imply such corrections for high-energy eigenstates. We will present a modified cosmic brane proposal that provides a universal explanation of such leading order corrections.
Our modified cosmic brane proposal is based on expanding the holographic state in a basis of fixed-area states where the areas of all extremal surfaces homologous to $R$ have small fluctuations Akers:2018fow ; Dong:2018seb ; Dong:2019piw . We review the decomposition of smooth holographic states into a fixed-area basis in Sec. 2.2, reformulating the cosmic brane proposal in this language. In the presence of two extremal surfaces $\gamma_{1}$ and $\gamma_{2}$, the cosmic brane proposal (ignoring bulk quantum corrections) takes the form
@@
To compare with the cosmic brane proposal (4), we rewrite (5) as
@@
We can now compare our proposal to the cosmic brane proposal. The key difference between the two proposals Eq. (4) and Eq. (6) is the order of optimization. For $n\geq 1$, we have two maximizations whose order can be swapped and thus the two proposals always agree. On the other hand, for $n<1$, the cosmic brane proposal is a minimax prescription whereas the modified cosmic brane proposal is a maximin prescription. Thus, in general, we can only conclude that $S_{n}^{MC}(R)\leq S_{n}^{C}(R)$, but they need not agree.
In Sec. 4, we provide sufficient conditions for an agreement between the two proposals. The original cosmic brane proposal considered two candidate saddles, and each saddle is smooth everywhere except that it has a cosmic brane at $\gamma_{i}$ which sources a conical defect of opening angle $\frac{2\pi}{n}$. We show that the two proposals agree whenever the cosmic brane sits at the minimal surface, i.e., when $A_{i}\leq A_{3-i}$ (evaluated in the corresponding saddle). On the other hand, it is possible that neither of the saddles satisfies this constraint. In such a case, the optimum in Eq. (6) may either be achieved by a subleading saddle in the cosmic brane proposal or at the entanglement phase transition boundary $A_{1}=A_{2}$, either of which results in leading order corrections to the cosmic brane proposal. As we explain in more detail later, the configurations at $A_{1}=A_{2}$ involve distributing the cosmic brane tension over both candidate RT surfaces, and these new configurations were not considered in the cosmic brane proposal.
To illustrate the presence of such corrections, we work out the example of $p\left(A_{1},A_{2}\right)$ being a Gaussian distribution in Sec. 5. In the simplest such setting, we demonstrate that the modified cosmic brane proposal agrees with the cosmic brane proposal for $n\geq 1$ as expected, whereas leading order corrections arise for $n<1$. The detailed analysis of Renyi entropies in an arbitrary Gaussian distribution is provided in Appendix A.
@@
2 Cosmic Brane Proposal
@@ 2.1 Review of the Proposal
Ref. Lewkowycz:2013nqa proposed that the dominant saddle $g_{n}$ respects the $\mathbb{Z}_{n}$ replica symmetry of the boundary path integral that cyclically permutes the $n$ boundary copies. This insight allows one to quotient the bulk geometry by this $\mathbb{Z}_{n}$ symmetry and provides an analytic continuation away from integer $n$, i.e., including the normalization factor we have
@@
where $\hat{g}_{n}$ is a solution to the equations of motion with a conical defect of opening angle $\frac{2\pi}{n}$ anchored to $\partial R$ in addition to satisfying the asymptotic boundary conditions defining the original state. [^1] The action $I\left[\hat{g}_{n}\right]$ excludes an explicit contribution from the conical defect Lewkowycz:2013nqa . Equivalently, this conical defect can be interpreted as arising from the insertion of a cosmic brane of tension $T_{n}=\frac{n-1}{4nG}$ Lewkowycz:2013nqa . With this understanding, Eq. (9) provides a natural analytic continuation to non-integer values of $n$ and defines the cosmic brane proposal for the holographic Renyi entropy for arbitrary $n$ at leading order in $G$, i.e.,
@@
It is important to note that the naive cosmic brane proposal is already subtle for $n<1$ in the presence of multiple extremal surfaces that serve as candidate RT surfaces for subregion $R$. When there are multiple extremal surfaces, at integer $n$ one naturally picks the solution with the least bulk action $I\left[\hat{g}_{n}\right]$ and it is natural to extend this rule to the analytic continuation. In the limit $n\to 1^{+}$, this picks out the minimal area surface, resulting in the RT formula.
However, if we instead considered the limit $n\to 1^{-}$, and if the naive cosmic brane proposal still chooses the solution with the least action $I\left[\hat{g}_{n}\right]$, it would pick out the maximal area surface among the two candidates, leading to a physically unreasonable answer. Thus, for $n<1$, it is natural to insist that the cosmic brane proposal pick out the solution with the largest action instead. With this rule, there is so far no obvious problem with the formula and we will take this to define the cosmic brane proposal in the presence of multiple extremal surfaces. Having said so, in this paper, we will demonstrate that even with this updated rule, the cosmic brane proposal can have corrections at leading order in $G$ in the presence of multiple extremal surfaces.
@@ 2.2 Reformulation in Terms of Fixed-Area States
Having discussed the cosmic brane proposal in terms of the gravitational path integral, we will now reformulate it in terms of fixed-area states. This will prove convenient later to compare with our modified proposal.
@@
We are now ready to reformulate the cosmic brane proposal in terms of fixed-area states.
@@
The cosmic brane proposal for the Renyi entropy is given by
@@
To see that Eq. (14) is equivalent to the formulation of the cosmic brane proposal reviewed in Sec. 2.1, we can use Eq. (13) to find the conditions for attaining the inner maximum in Eq. (14): e.g. for $i=1$, we have
@@
The action of the fixed-area saddle can be divided into two parts: a localized contribution from the surfaces $\gamma_{1,2}$ and the remaining contribution from the rest of the spacetime away from the surfaces Dong:2018seb ; i.e,
@@
where we remind the reader that $\phi_{i}$ is the opening angle at surface $\gamma_{i}$ and thus, the localized contributions arise from Ricci scalar delta functions due to the presence of conical defects at $\gamma_{1,2}$. To take the derivative with respect to $A_{i}$, we follow the argument of Ref. Dong:2018seb and note that differentiating $I_{away}\left[g_{A_{1},A_{2}}\right]$ with respect to $\phi_{i}$ gives $-A_{i}/8\pi G$, and thus its Legendre transform $I\left[g_{A_{1},A_{2}}\right]$ satisfies
@@
Using this, Eq. (15) and Eq. (16) imply that $\phi_{1}=\frac{2\pi}{n}$ and $\phi_{2}=2\pi$. Again, by symmetry, we have $\phi_{1}=2\pi$ and $\phi_{2}=\frac{2\pi}{n}$ for $i=2$. These are precisely the candidate saddles that one obtains from the cosmic brane proposal, with Eq. (15) indicating the presence of a cosmic brane of tension $T_{n}$ at $\gamma_{i}$, as discussed in Sec. 2.1. Finally, as discussed in Sec. 2.1, the maximization (minimization) over $i$ for $n\geq 1$ ($n<1$) results in the optimum configuration chosen in the cosmic brane proposal.
While we showed Eq. (14) to be equivalent to the cosmic brane proposal for holographic states prepared by a smooth gravitational path integral, Eq. (14) can be taken as the definition of the cosmic brane proposal for more general probability distributions over fixed-area states.
@@ 3 Modified Cosmic Brane Proposal
We can now discuss the effect of bulk quantum corrections to our modified cosmic brane proposal. In general, accounting for bulk quantum corrections is complicated since the replica-symmetry breaking contributions need to be carefully summed over in order to analytically continue in $n$. We will discuss this further in Sec. 7.
@@ 4 Comparing the Two Proposals
For $n\geq 1$, we have two maximizations whose order can always be swapped. Thus, the two proposals agree in this case. For $n<1$, the cosmic brane proposal is a minimax prescription whereas the modified cosmic brane proposal is a maximin prescription. From the well-known max-min inequality (see for instance Ref. boyd2004convex ), we obtain the required inequality. ∎
We will now find sufficient conditions for the two proposals to agree for $n<1$. The cosmic brane proposal has two candidate saddles, each with a cosmic brane sitting at $\gamma_{i}$. We will find that the two proposals agree as long as at least one of these saddles satisfies the constraint that $\gamma_{i}$ is the minimal surface among $\gamma_{1,2}$ (which we will refer to as the minimality constraint).
@@
The cosmic brane proposal for $n<1$ is given by
@@
We will now determine sufficient conditions under which these proposals agree.
@@
Thus, we see that the two proposals agree as long as at least one of the candidate saddles in the cosmic brane proposal satisfies the minimality constraint. This implies that the two proposals can only disagree when neither of these saddles satisfies the minimality constraint. In such a situation, the modified cosmic brane proposal will pick either a subleading saddle (for the cosmic brane proposal) that satisfies the minimality constraint, or a diagonal saddle $\boldsymbol{\mathbf{A^{(D)}}}$ that satisfies $A_{1}=A_{2}$.555 Related boundary value dominance has shown up in similar analyses of logarithmic negativity 2022PhRvL.129f1602V ; 2021arXiv211200020V . We will explore various controlled examples in Sec. 5 and Sec. 6 where $\boldsymbol{\mathbf{A^{(D)}}}$ does indeed dominate for $n<1$, resulting in leading order corrections to the cosmic brane proposal. We expect this to be a generic feature for $n<1$.
Before moving on, it is illuminating to discuss the geometry of the saddle $\boldsymbol{\mathbf{A^{(D)}}}$ to contrast with the cosmic brane proposal in situations where the state is defined by a smooth gravitational path integral. On the diagonal, we can write $\min\left[A_{1},A_{2}\right]=xA_{1}+(1-x)A_{2}$ for any $x$. Then, the maximizing conditions for the modified cosmic brane proposal are
@@
where we look for solutions that satisfy the constraint $A_{1}=A_{2}$, thus providing enough conditions to solve for $x$. The conditions Eq. (35) and Eq. (36) can be interpreted as introducing cosmic branes with distributed tensions $T_{n,1}=xT_{n}$ and $T_{n,2}=(1-x)T_{n}$ at the surfaces $\gamma_{1}$ and $\gamma_{2}$ respectively. This is in contrast with the cosmic brane proposal where the cosmic brane of tension $T_{n}$ lies at a single surface.

### v3

@@ Abstract
We propose a new formula for computing holographic Renyi entropies in the presence of multiple extremal surfaces. Our proposal is based on computing the wave function in the basis of fixed-area states and assuming a diagonal approximation for the Renyi entropy. For Renyi index $n\geq 1$, our proposal agrees with the existing cosmic brane proposal for holographic Renyi entropy. For $n<1$, however, our modified cosmic brane proposal predicts a new phase with leading order (in Newton’s constant $G$) corrections to the original cosmic brane proposal, even far from entanglement phase transitions and when bulk quantum corrections are unimportant. Recast in terms of optimization over fixed-area states, the difference between the two proposals can be understood to come from the order of optimization: for $n<1$, the original cosmic brane proposal is a minimax prescription whereas our proposal is a maximin prescription. We demonstrate the presence of such leading order corrections using illustrative examples. In particular, our proposal reproduces existing results in the literature for the PSSY model and high-energy eigenstates, providing a universal explanation for previously found leading order corrections to the $n<1$ Renyi entropies.
@@ 1 Introduction
The Renyi entropies at integer $n$ can be computed using the replica trick in the boundary CFT Calabrese:2004eu . The insight of Ref. Lewkowycz:2013nqa , as we shall review in Sec. 2.1, was to propose that the corresponding dominant bulk gravitational saddle is replica symmetric. Then, quotienting the bulk geometry by the replica symmetry, one obtains a geometry with an additional conical defect of opening angle $\frac{2\pi}{n}$ anchored to $\partial R$. Such a conical defect could equivalently be interpreted as being induced by the insertion of a “cosmic brane” of appropriate tension Dong:2016fnf ; Lewkowycz:2013nqa . This cosmic brane proposal provides a natural continuation away from integer $n$, and the RT formula follows from it in the limit $n\to 1$. Moreover, it was shown in Ref. Dong:2016fnf that the refined Renyi entropy is then computed by the area of the cosmic brane in this new spacetime, providing a generalization of the RT formula. Henceforth, we will refer to either of the above proposals for Renyi entropy and refined Renyi entropy as the original cosmic brane proposal.
In this paper, we will demonstrate that while the original cosmic brane proposal is correct in many situations, it can fail even at leading order (in $G$) in large regions of parameter space. In particular, we will show that such corrections appear quite generically for Renyi index $n<1$ in the presence of multiple extremal surfaces. Such corrections have previously been noticed by Refs. Dong:2021oad ; Akers:2022max in the PSSY model of black hole evaporation Penington:2019kki . Moreover, the results of Ref. Murthy:2019qvb interpreted in a holographic context also imply such corrections for high-energy eigenstates. We will present a modified cosmic brane proposal that provides a universal explanation of such leading order corrections.
Our modified cosmic brane proposal is based on expanding the holographic state in a basis of fixed-area states where the areas of all extremal surfaces homologous to $R$ have small fluctuations Akers:2018fow ; Dong:2018seb ; Dong:2019piw . We review the decomposition of smooth holographic states into a fixed-area basis in Sec. 2.2, reformulating the original cosmic brane proposal in this language. In the presence of two extremal surfaces $\gamma_{1}$ and $\gamma_{2}$, the original cosmic brane proposal (ignoring bulk quantum corrections) takes the form
@@
To compare with the original cosmic brane proposal (4), we rewrite (5) as
@@
We can now compare our proposal to the original cosmic brane proposal. The key difference between the two proposals Eq. (4) and Eq. (6) is the order of optimization. For $n\geq 1$, we have two maximizations whose order can be swapped and thus the two proposals always agree. On the other hand, for $n<1$, the original cosmic brane proposal is a minimax prescription whereas the modified cosmic brane proposal is a maximin prescription. Thus, in general, we can only conclude that $S_{n}^{MC}(R)\leq S_{n}^{C}(R)$, but they need not agree.
In Sec. 4, we provide necessary and sufficient conditions for an agreement between the two proposals. The original cosmic brane proposal considered two candidate saddles, and each saddle is smooth everywhere except that it has a cosmic brane at $\gamma_{i}$ which sources a conical defect of opening angle $\frac{2\pi}{n}$. We show that the two proposals agree whenever the cosmic brane sits at the minimal surface, i.e., when $A_{i}\leq A_{3-i}$ (evaluated in the corresponding saddle). On the other hand, it is possible that neither of the saddles satisfies this constraint. In such a case, the optimum in Eq. (6) may either be achieved by a subleading saddle in the original cosmic brane proposal or at the entanglement phase transition boundary $A_{1}=A_{2}$, either of which results in leading order corrections to the original cosmic brane proposal. As we explain in more detail later, the configurations at $A_{1}=A_{2}$ involve distributing the cosmic brane tension over both candidate RT surfaces, and these new configurations were not considered in the original cosmic brane proposal.
To illustrate the presence of such corrections, we work out the example of $p\left(A_{1},A_{2}\right)$ being a Gaussian distribution in Sec. 5. In the simplest such setting, we demonstrate that the modified cosmic brane proposal agrees with the original cosmic brane proposal for $n\geq 1$ as expected, whereas leading order corrections arise for $n<1$. The detailed analysis of Renyi entropies in an arbitrary Gaussian distribution is provided in Appendix A.
@@
2 Original Cosmic Brane Proposal
@@ 2.1 Review of the Proposal
Ref. Lewkowycz:2013nqa proposed that the dominant saddle $g_{n}$ respects the $\mathbb{Z}_{n}$ replica symmetry of the boundary path integral that cyclically permutes the $n$ boundary copies. This insight allows one to quotient the bulk geometry by this $\mathbb{Z}_{n}$ symmetry and provides a natural continuation away from integer $n$, i.e., including the normalization factor we have
@@
where $\hat{g}_{n}$ is a solution to the equations of motion with a conical defect of opening angle $\frac{2\pi}{n}$ anchored to $\partial R$ in addition to satisfying the asymptotic boundary conditions defining the original state. [^1] The action $I\left[\hat{g}_{n}\right]$ excludes an explicit contribution from the conical defect Lewkowycz:2013nqa . Equivalently, this conical defect can be interpreted as arising from the insertion of a cosmic brane of tension $T_{n}=\frac{n-1}{4nG}$ Lewkowycz:2013nqa . With this understanding, Eq. (9) provides a natural continuation to non-integer values of $n$ [^2] Note that the continuation need not be analytic since there may generally be phase transitions in the $G\to 0$ limit, in which case one cannot apply Carlson’s theorem as one could for the boundary CFT at finite central charge. and defines the original cosmic brane proposal for the holographic Renyi entropy for arbitrary $n$ at leading order in $G$, i.e.,
@@
It is important to note that the naive version of the original cosmic brane proposal is already subtle for $n<1$ in the presence of multiple extremal surfaces that serve as candidate RT surfaces for subregion $R$. When there are multiple extremal surfaces, at integer $n$ one naturally picks the solution with the least bulk action $I\left[\hat{g}_{n}\right]$ and it is natural to extend this rule to the continuation. In the limit $n\to 1^{+}$, this picks out the minimal area surface, resulting in the RT formula.
However, if we instead considered the limit $n\to 1^{-}$, and if the original cosmic brane proposal still chooses the solution with the least action $I\left[\hat{g}_{n}\right]$, it would pick out the maximal area surface among the two candidates, leading to a physically unreasonable answer. Thus, for $n<1$, it is natural to insist that the original cosmic brane proposal pick out the solution with the largest action instead. With this rule, there is so far no obvious problem with the formula and we will take this to define the original cosmic brane proposal in the presence of multiple extremal surfaces. Having said so, in this paper, we will demonstrate that even with this updated rule, the original cosmic brane proposal can have corrections at leading order in $G$ in the presence of multiple extremal surfaces.
@@ 2.2 Reformulation in Terms of Fixed-Area States
Having discussed the original cosmic brane proposal in terms of the gravitational path integral, we will now reformulate it in terms of fixed-area states. This will prove convenient later to compare with our modified proposal.
@@
We are now ready to reformulate the original cosmic brane proposal in terms of fixed-area states.
@@
The original cosmic brane proposal for the Renyi entropy is given by
@@
To see that Eq. (14) is equivalent to the formulation of the original cosmic brane proposal reviewed in Sec. 2.1, we can use Eq. (13) to find the conditions for attaining the inner maximum in Eq. (14): e.g. for $i=1$, we have
@@
In general, the action of the fixed-area saddle can be divided into two parts: a localized contribution from the surfaces $\gamma_{1,2}$ and the remaining contribution from the rest of the spacetime away from the surfaces Dong:2018seb ; i.e,
@@
where we remind the reader that $\phi_{i}$ is the opening angle at surface $\gamma_{i}$ and thus, the localized contributions arise from Ricci scalar delta functions due to the presence of conical defects at $\gamma_{1,2}$. To take the derivative with respect to $A_{i}$ for a smooth holographic state, we follow the argument of Ref. Dong:2018seb and note that differentiating $I_{away}\left[g_{A_{1},A_{2}}\right]$ with respect to $\phi_{i}$ gives $-A_{i}/8\pi G$, and thus its Legendre transform $I\left[g_{A_{1},A_{2}}\right]$ satisfies
@@
Using this, Eq. (15) and Eq. (16) imply that $\phi_{1}=\frac{2\pi}{n}$ and $\phi_{2}=2\pi$. Again, by symmetry, we have $\phi_{1}=2\pi$ and $\phi_{2}=\frac{2\pi}{n}$ for $i=2$. These are precisely the candidate saddles that one obtains from the original cosmic brane proposal, with Eq. (15) indicating the presence of a cosmic brane of tension $T_{n}$ at $\gamma_{i}$, as discussed in Sec. 2.1. Finally, as discussed in Sec. 2.1, the maximization (minimization) over $i$ for $n\geq 1$ ($n<1$) results in the optimum configuration chosen in the original cosmic brane proposal.
While we showed Eq. (14) to be equivalent to the original cosmic brane proposal for holographic states prepared by a smooth gravitational path integral, Eq. (14) can be taken as the definition of the original cosmic brane proposal for more general probability distributions over fixed-area states. In fact, this is natural since the effect of inserting a cosmic brane of tension $T_{n}$ is captured by Eq. (15) and Eq. (16) quite generally, even for non-smooth holographic states like fixed-area states.
@@ 3 Modified Cosmic Brane Proposal
We can now discuss the effect of bulk quantum corrections to our modified cosmic brane proposal. In general, accounting for bulk quantum corrections is complicated since the replica-symmetry breaking contributions need to be carefully summed over in order to continue to non-integer $n$. We will discuss this further in Sec. 7.
@@ 4 Comparing the Two Proposals
For $n\geq 1$, we have two maximizations whose order can always be swapped. Thus, the two proposals agree in this case. For $n<1$, the original cosmic brane proposal is a minimax prescription whereas the modified cosmic brane proposal is a maximin prescription. From the well-known max-min inequality (see for instance Ref. boyd2004convex ), we obtain the required inequality. ∎
We will now find necessary and sufficient conditions for the two proposals to agree for $n<1$. The original cosmic brane proposal has two candidate saddles, each with a cosmic brane sitting at $\gamma_{i}$. We will find that the two proposals agree if and only if at least one of these saddles satisfies the constraint that $\gamma_{i}$ is the minimal surface among $\gamma_{1,2}$ (which we will refer to as the minimality constraint).
@@
The original cosmic brane proposal for $n<1$ is given by
@@
We will now determine necessary and sufficient conditions under which these proposals agree.
@@
Thus, we see that the two proposals agree as long as at least one of the candidate saddles in the original cosmic brane proposal satisfies the minimality constraint. This implies that the two proposals can only disagree when neither of these saddles satisfies the minimality constraint. In such a situation, the modified cosmic brane proposal will pick either a subleading saddle (for the original cosmic brane proposal) that satisfies the minimality constraint, or a diagonal saddle $\boldsymbol{\mathbf{A^{(D)}}}$ that satisfies $A_{1}=A_{2}$.666 Related boundary value dominance has shown up in similar analyses of logarithmic negativity 2022PhRvL.129f1602V ; 2021arXiv211200020V . We will explore various controlled examples in Sec. 5 and Sec. 6 where $\boldsymbol{\mathbf{A^{(D)}}}$ does indeed dominate for $n<1$, resulting in leading order corrections to the original cosmic brane proposal. We expect this to be a generic feature for $n<1$.
Before moving on, it is illuminating to discuss the geometry of the saddle $\boldsymbol{\mathbf{A^{(D)}}}$ to contrast with the original cosmic brane proposal in situations where the state is defined by a smooth gravitational path integral. On the diagonal, we can write $\min\left[A_{1},A_{2}\right]=xA_{1}+(1-x)A_{2}$ for any $x$. Then, the maximizing conditions for the modified cosmic brane proposal are
@@
where we look for solutions that satisfy the constraint $A_{1}=A_{2}$, thus providing enough conditions to solve for $x$. The conditions Eq. (35) and Eq. (36) can be interpreted as introducing cosmic branes with distributed tensions $T_{n,1}=xT_{n}$ and $T_{n,2}=(1-x)T_{n}$ at the surfaces $\gamma_{1}$ and $\gamma_{2}$ respectively. This is in contrast with the original cosmic brane proposal where the cosmic brane of tension $T_{n}$ lies at a single surface.